# The InSAR absolute phase amid singularities

### v1

@@ 2.2.2 Evaluation
In the language of exterior calculus [21], the mapping $\gamma$ determines $d\varphi_{W^{+}}=\gamma^{\star}(d\varphi_{\mathbb{C}^{+}})$ as the pullback from $d\varphi_{\mathbb{C}^{+}}$.
The evaluation is illustrated in Fig. 2d) for the path $\Gamma$ from $\mathbf{w}_{1}$ to $\mathbf{w}^{\prime}$, shown in red. Mapping this path into $\mathbb{C}^{+}$, the path $\gamma\circ\Gamma$ is obtained. The unwrapped phase $\Delta\phi(\Gamma)$ is the angle spanned by $\gamma\circ\Gamma$ as seen from the origin [8].
2.2.3 Path dependence of phase unwrapping
For fixed endpoints, the unwrapped phase difference $\Delta\phi(\Gamma)$ can depend on the path $\Gamma$. In Fig. 2b), paths $\Gamma$ and $\tilde{\Gamma}$ both start at $w_{1}$ and end at $w^{\prime}$. In the complex plane (Fig. 2d), $\gamma\circ\Gamma$ and $\gamma\circ\tilde{\Gamma}$ also have identical endpoints. However, the unwrapped phases $\Delta\phi(\Gamma)$ and $\Delta\phi(\tilde{\Gamma})$ are not equal, indicating path dependence. The difference of $2\pi$ is due to the paths passing the origin on opposite sides. The path dependence ultimately derives from the phase singularity in $W$. If we were to first follow $\Gamma$ and then $\tilde{\Gamma}$ in the opposite direction (Fig. 2b), we would encircle the phase singularity in a loop, leading to a total unwrapped phase around the loop of $-2\pi$.
Any loop yields an unwrapped phase difference of $2\pi N$, where $N\in\mathbb{N}$ [8]. The winding number $N$ is an integer that describes the number of phase cycles along the loop, or equivalently, how many times the loop mapped into the complex plane wraps around the origin. Alternative names for $N$ include residue (in InSAR), topological charge or topological quantum number (in physics).
@@ 3.1 General definition
The absolute phase is unwrapped along a pre-specified absolute path $\Gamma^{\mathrm{a}}_{\mathbf{w}_{2}}\in W$ that ends at $\mathbf{w}_{2}$, representing the configuration at the secondary acquisition. The natural path $\Gamma_{a}$ handles the path dependence of phase unwrapping by restricting attention to just one path.
@@
[…]
@@ 3.2 Absolute phase singularities
I define an absolute phase singularity to be any point $\mathbf{w}_{2}$ for which the coherence vanishes somewhere along the path $\Gamma^{\mathrm{a}}_{\mathbf{w}_{2}}$. The set of all such points, $A$, also contains the phase singularities. The absolute phase is defined for all $\mathbf{w}_{2}$ in
@@
In Fig. 2, the absolute phase singularities $A$ are located on a ray that extends from the phase singularity at the origin. For any $\mathbf{w}\in A$, the absolute paths, namely the straight line paths from $\mathbf{x}_{1}$, intersect the phase singularity at the origin. Across this phase singularity, the absolute phase jumps by $2\pi$. This absolute phase singularity is thus similar to a branch cut of the complex logarithm.
@@ 3.3 Interpretation
The absolute phase $\phi^{\mathrm{a}}(\mathbf{w}_{1},\mathbf{w}_{1})=0$ for identical primary and secondary configuration parameters, by definition.
The absolute phase $\phi^{\mathrm{a}}(\mathbf{w}_{2},\mathbf{w}_{1})$ is congruent [8, 7] with the wrapped phase $\phi(\mathbf{w}_{2},\mathbf{w}_{1})$ in that
$\arg\exp(\mathsf{j}\phi^{\mathrm{a}}(\mathbf{w}_{2},\mathbf{w}_{1}))=\varphi(\mathbf{w}_{2})\,.$ (14)
The two are equivalent modulo $2\pi$, as $\varphi(\mathbf{w}_{2},\mathbf{w}_{1})=[\phi^{\mathrm{a}}{\mathbf{w}_{2},\mathbf{w}_{1}}]$. While the wrapped phase changes sign when the primary and secondary acquisition are interchanged, $\varphi(\mathbf{w}_{2},\mathbf{w}_{1})=-\varphi(\mathbf{w}_{1},\mathbf{w}_{2})$, there is no such guarantee for the absolute phase. Assuming both $\phi^{\mathrm{a}}{\mathbf{w}_{2},\mathbf{w}_{1}}$ and $\phi^{\mathrm{a}}{\mathbf{w}_{1},\mathbf{w}_{2}}$ are defined, (14) implies that $[\phi^{\mathrm{a}}{\mathbf{w}_{2},\mathbf{w}_{1}}]=[-\phi^{\mathrm{a}}{\mathbf{w}_{1},\mathbf{w}_{2}}]$, but the integer multiple of $2\pi$ will depend on the coherence mapping and the path $\Gamma^{\mathrm{a}}$.
@@
is the default. It connects $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ through a straight line. It meets both desiderata for $\Gamma^{\mathrm{a}}$: i) independence of $\gamma(\mathbf{w})$, and ii) smooth dependence on the endpoint $x_{2}$. Its definition makes sense for star-shaped $W$ (with center $\mathbf{w}_{1}$). For more general $W$, different choices of $\Gamma^{\mathrm{a}}$ are required. Points of discontinuity are then included in the absolute phase singularities $A$, implying that the absolute phase is not defined there and rescuing continuity of the absolute phase in $W\setminus A$. The arbitrary selection of an absolute path impinges on the absolute phase, reinforcing that the wrapped phase is the fundamental physical quantity in interferometry.
@@ 3.4 Analogy
A familiar analogy unrelated to interferometry illustrates these concepts: The offset in the time of day on a spherical Earth relative to a standard time. The time zones are a discretized variant of this offset [9]. The wrapped phase for any point on the Earth’s surface corresponds to the difference between local solar time and a reference time, with a local solar time of zero when the sun is lowest below the horizon. (For simplicity, I will neglect the tilt of the Earth’s axis and the eccentricity of its orbit, thus equating mean and apparent solar time.) The sine of the amplitude of the sun’s elevation angle relative to the local horizon may be associated with a coherence magnitude, with the Earth’s poles as phase singularities [9].
An absolute phase is induced by definition of an absolute path with an initial point of 0 latitude and 0 longitude. Define the absolute path to be the shortest among all paths that first run along the equator and then along a line of longitude, choosing to go East first for points on the 180 meridian. The absolute phase in this case is restricted to the interval $(-\pi,\pi]$ and establishes the notion of a calendar day for the entire Earth. The jump by 1 day in time (or $2\pi$ in phase) identifies the 180 longitude line as a simplified international date line, or an absolute phase singularity based on the convention from the previous paragraph. A different absolute path – for instance, only going west along the equator – would induce a different absolute phase or calendar day. More generally, the unwrapped phase along a path corresponds to the time difference between local solar time and the time shown by a clock of a traveler who has traversed the path and who initially synchronized their clock. A traveler going around the world in eighty days in an eastward loop perceives a time difference of one day [22].
@@ 4.1 Definition
In an ideal case, only time distinguishes the secondary from the primary observational configuration. Keeping the primary acquisition at time $t_{1}$ fixed, we can let $W\cong\mathbb{R}$ and $\mathbf{w}\equiv t$. From (12) and (15), the differential InSAR absolute phase at a secondary time $t_{2}$ is defined by
[…]
@@
The coherence $\gamma(t)=\xi(t)+\mathsf{j}\eta(t)$ is evaluated for a fixed primary at $t_{1}$ and variable intermediate secondary acquisition at $t$. It can be convenient to use a dimensionless time $s=(t-t_{1})/(t_{2}-t_{1})$, in which case the limits are 0 and 111. Then, $\phi^{\mathrm{a}}(s)$ for $0\leq s<1$ is the absolute phase for “intermediate” secondary acquisitions prior to the “final” secondary acquisition at $s=1$.
@@ 4.2 Modeling framework
[…]
where the differential one-form $d\varphi_{B^{+}}$ is analogous to $d\varphi_{W^{+}}$. This simplified framework is mathematically convenient because $B$ now has the same role as $W$. In particular, we call $\mathbf{b}_{2}$ an absolute phase singularity if the phase is undefined somewhere along $\Gamma^{\mathrm{a}}_{\mathbf{b}_{2}}$. The domain of $\phi^{\mathrm{a}}_{B^{+}}$ is denoted $B^{\mathrm{a}}$. As for $\phi^{\mathrm{a}}$, we will adopt the convex combination absolute path (15). For instance, if $\mathbf{b}$ encodes scatterer positions, the scatterers will move at a uniform rate from $\mathbf{b}_{1}$ to $\mathbf{b}_{2}$.
@@ 4.3.1 Point scatterer
For a point-like scatterer, the general definition of the absolute phase (17) is proportional to the difference in optical path length, as expected [1]. Assuming a monostatic configuration, zero spatial baseline, the plane wave approximation and no atmospheric or other phase screens, the single-look coherence
$\gamma_{\mathrm{point}}(t)=\exp\left(\mathsf{j}2k\left[d(t)-d(t_{1})\right]\right)\,,$ (19)
where $k$ is the scalar wavenumber, $d$ the scatterer’s line-of-sight distance, and the primary acquisition at $t_{1}$ is fixed. Multilook observations of a non-decorrelating, uniformly moving surface are also described by (19)–(20). Previously [1, 2], the absolute phase of such scatterers was taken as
$\phi^{\mathrm{a}}_{\mathrm{point}}(t_{2};t_{1})=2k\left[d(t_{2})-d(t_{1})\right]\,.$ (20)
@@
$d\varphi_{t}(\mathrm{d}t)=2k\frac{\mathrm{d}d}{\mathrm{d}t}\,\mathrm{d}t=2k\,\mathrm{d}d\,$ (21)
is an exact form, so the integral only depends on the range difference $d(t_{2})-d(t_{1})$ and not on the path $d(t)$.
@@ 4.3.2 Three surface segments
The three surface segments from the introduction differ from a single surface or point scatterer in that a cyclic change can induce a non-zero absolute phase. To model the system in the simplified framework, let the state $\mathbf{b}=[b_{1},b_{2},b_{3}]^{T}$ encode the position of the three segments. Assuming the returns from different segments are uncorrelated and of equal strength [23], let $b_{i}(s)$ be the position of surface $i$ at dimensionless time $s$, nondimensionalized such that the modeled population coherence is
$\gamma(s)=\frac{1}{3}\sum_{i=1}^{3}\exp^{-\mathsf{j}\left(w_{i}(s)-w_{i}(0)\right)}\,.$ (22)
@@
Each segment starts out at $b_{i}(0)=0$ for the fixed primary acquisition at $s=0$ and acts like a non-decorrelating point-like scatterer from (19). The temporal evolution shown in Fig. 3a is cyclic, as all segments return to their initial position. Our goal is to analyze the nonzero value of the absolute phase $\phi^{\mathrm{a}}(1)=-2\pi$ for the loop that we found in Fig. 1.
@@
Figure 3: Differential InSAR model of three surface segments with positions $\mathbf{b}=[b_{1},b_{2},b_{3}]^{T}$. a) The purple loop with basepoint $\mathbf{b}_{1}=[0,0,0]$ (white cross) in parameter space describes the temporal evolution from Fig. 1, it encircles a phase singularity (black line) in the plane $b_{1}=0$. b–d) Wrapped phase (b), coherence (c) and absolute phase in the modeling framework of Sec. 4.2 (d) shown in the plane $b_{1}=0$ for fixed primary $\mathbf{b}_{1}$, along with phase singularities (black dots) and absolute phase singularities (gray lines).
@@
The nonzero absolute phase around the loop relates to the loop enclosing a phase singularity, one of the two dark lines in Fig. 3a). In the plane $b_{1}=0$, the wrapped phase $\phi(\mathbf{b})$ for fixed primary $\mathbf{b}_{1}=0$ in Fig. 3b spirals around the phase singularity, such that the unwrapped phase around the loop gives the absolute phase $\phi^{\mathrm{a}}(1)=-2\pi$. As $\rho>0$ along the loop (Fig. 3c), this absolute phase is well defined.
@@
The phase singularities induce absolute phases singularities. Rather than considering the loop, we apply the convention of Sec. 4.2, restricting trajectories to the line segment from $\mathbf{b}_{1}=0$ to secondary $\mathbf{b}$. The absolute phase in parameter space, $\phi^{\mathrm{a}}_{B}$ is shown in Fig. 3d). The absolute phase singularities $A\in B$ in the plane $b_{1}=0$ are rays extending from each phase singularity, across which the absolute phase jumps by $2\pi$. They leave no trace in the wrapped phase. In Sec. 6.1, it will be shown that the discontinuous jumps arise from the path-dependence of the phase one-form illustrated by the loop encircling the phase singularity.

### v3

@@ 2.2.3 Path dependence of phase unwrapping
Any loop yields an unwrapped phase difference of $2\pi N$, where $N\in\mathbb{N}$ [fulton97]. The winding number $N$ is an integer that describes the number of phase cycles along the loop, or equivalently, how many times the loop mapped into the complex plane wraps around the origin. Alternative names for $N$ include residue (in InSAR), topological charge or topological quantum number (in physics).
@@ 3.1 General definition
The absolute phase is unwrapped along a pre-specified absolute path $\Gamma^{\mathrm{a}}_{\mathbf{w}_{2}}\in W$ that ends at $\mathbf{w}_{2}$, representing the configuration at the secondary acquisition. The natural path $\Gamma_{a}$ handles path dependence of phase unwrapping by restricting attention to just one path. Its specification requires a choice and will depend on context. We will assume a reasonable choice has been made, deferring an in-depth discussion to later.
@@
$\phi^{\mathrm{a}}(\mathbf{w}_{2})\equiv\Delta\phi(\Gamma^{\mathrm{a}}_{\mathbf{w}_{2}})=\int_{\Gamma^{\mathrm{a}}_{\mathbf{w}_{2}}}d\varphi_{W^{+}}\,.$ (12)
@@ 3.2 Absolute phase singularities
I define an absolute phase singularity to be any point $\mathbf{w}_{2}$ for which the coherence vanishes somewhere along the path $\Gamma^{\mathrm{a}}_{\mathbf{w}_{2}}$. The set of all such points, $A$, contains the phase singularities (for which the wrapped phase is undefined) as well as points for which the wrapped phase is defined. The absolute phase is not defined in $A$, but it is for all $\mathbf{w}_{2}$ in
@@
In Fig. 2e, the absolute phase singularities $A$ lie on a ray extending from the phase singularity away from $\mathbf{w}1$. For any $\mathbf{w}\in A$, the absolute paths, namely the straight line paths from $\mathbf{w}_{1}$, intersect the phase singularity. Across this phase singularity, the absolute phase jumps by $2\pi$. This absolute phase singularity is thus similar to a branch cut of the complex logarithm.
@@ 3.3 Interpretation
The absolute phase $\phi^{\mathrm{a}}(\mathbf{w}_{1},\mathbf{w}_{1})=0$ for identical primary and secondary configuration parameters, by definition, assuming that the absolute path $\Gamma^{\mathrm{a}}_{\mathbf{w}_{1}}$ is a constant path.
The absolute phase $\phi^{\mathrm{a}}(\mathbf{w}_{2},\mathbf{w}_{1})$ is congruent [fulton97, ghiglia98] with the wrapped phase $\varphi(\mathbf{w}_{2},\mathbf{w}_{1})$ in that
$\arg\exp(\mathsf{j}\phi^{\mathrm{a}}(\mathbf{w}_{2},\mathbf{w}_{1}))=\varphi(\mathbf{w}_{2},\mathbf{w}_{1})\,.$ (14)
The two are equivalent modulo $2\pi$, i.e., $\varphi(\mathbf{w}_{2},\mathbf{w}_{1})=[\phi^{\mathrm{a}}(\mathbf{w}_{2},\mathbf{w}_{1})]$. While the wrapped phase changes sign when the primary and secondary acquisition are interchanged, $\varphi(\mathbf{w}_{2},\mathbf{w}_{1})=-\varphi(\mathbf{w}_{1},\mathbf{w}_{2})$, there is no such guarantee for the absolute phase. Assuming both $\phi^{\mathrm{a}}(\mathbf{w}_{2},\mathbf{w}_{1})$ and $\phi^{\mathrm{a}}(\mathbf{w}_{1},\mathbf{w}_{2})$ are defined, (14) implies that $\phi^{\mathrm{a}}(\mathbf{w}_{2},\mathbf{w}_{1})]=[-\phi^{\mathrm{a}}(\mathbf{w}_{1},\mathbf{w}_{2})]$, but the integer multiple of $2\pi$ will depend on the coherence mapping and the path $\Gamma^{\mathrm{a}}$.
@@
is the default. It connects $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ through a straight line. It meets all three desiderata for $\Gamma^{\mathrm{a}}$: i) constancy of $\Gamma^{\mathrm{a}}_{\mathbf{w}_{1}}$; ii) independence of $\gamma(\mathbf{w})$, and iii) smooth dependence on the endpoint $x_{2}$. Its definition makes sense for star-shaped $W$ (with center $\mathbf{w}_{1}$). For more general $W$, different choices of $\Gamma^{\mathrm{a}}$ are required. If the absolute path has discontinuities, these are then included in the absolute phase singularities $A$, implying that the absolute phase is not defined on such points and rescuing continuity of the absolute phase in $W\setminus A$.
@@ 3.4 Analogy
A familiar analogy unrelated to interferometry illustrates these concepts: The offset in the time of day on a spherical Earth relative to a standard time. The time zones are a discretized variant of this offset [dennis01]. The wrapped phase for any point on the Earth’s surface corresponds to the difference between local solar time and a reference time, with a local solar time of zero when the sun is lowest below the horizon. (For simplicity, I will neglect the tilt of the Earth’s axis and the eccentricity of its orbit, thus equating mean and apparent solar time.) The sine of the amplitude of the sun’s elevation angle relative to the local horizon may be associated with a coherence magnitude, with the Earth’s poles as phase singularities [dennis01].
An absolute phase is induced by definition of an absolute path with an initial point of 0 latitude and 0 longitude. Define the absolute path to be the shortest among all paths that first run along the equator and then along a line of longitude, choosing to go East first for points on the 180 meridian. The absolute phase in this case is restricted to the interval $(-\pi,\pi]$ and establishes the notion of a calendar day for the entire Earth. The jump by 1 day in time (or $2\pi$ in phase) identifies the 180 longitude line as a simplified international date line, or an absolute phase singularity based on the convention from the previous subsection. A different absolute path – for instance, only going west along the equator – would induce a different absolute phase or calendar day. More generally, the unwrapped phase along a path corresponds to the time difference between local solar time and the time shown by the clock of a traveler who has traversed the path and who initially synchronized their clock. A traveler going around the world in eighty days in an eastward loop perceives a time difference of one day [verne73].
@@ 4.1 Definition
In an ideal case, only time distinguishes the secondary from the primary configuration. Keeping the primary acquisition at time $t_{1}$ fixed, we can let $W\cong\mathbb{R}$ and $\mathbf{w}\equiv t$. From (12) and (15), the differential InSAR absolute phase at a secondary time $t_{2}$ is defined by
$\phi^{\mathrm{a}}(t_{2})$ $=\int_{t_{1}}^{t_{2}}d\varphi_{t}$ (16)
@@
The coherence $\gamma(t)=\xi(t)+\mathsf{j}\eta(t)$ is evaluated for a fixed primary at $t_{1}$ and variable intermediate secondary acquisition at $t$. It can be convenient to use a dimensionless time $s=(t-t_{1})/(t_{2}-t_{1})$, in which case the limits are 0 and $1$. Then, $\phi^{\mathrm{a}}(s)$ for $0\leq s<1$ is the absolute phase for “intermediate” secondary acquisitions prior to the “final” secondary acquisition at $s=1$.
In differential InSAR, the absolute path basically requires no choice. The observational definition of the absolute phase in (17) is the unique definition of a phase-like quantity in $\mathbb{R}$ that fulfills three intuitive axioms, for fixed $t_{1}$ and continuous $\gamma(t_{2})$. Axiom 1: it depends continuously on secondary time $t_{2}$; axiom 2: it is congruent with the wrapped phase $\varphi(t_{2})$ for all $t_{2}$; and axiom 3: it is zero for $t_{1}=t_{2}$. The statement then follows from Proposition 1.34 in [hatcher01]; in the theory of covering spaces, it is an application of the unique lifting property of a mapping into $S^{1}$. Axioms 1 and 2 determine the phase-like quantity up to an integer multiple of $2\pi$, while axiom 3 fixes this integer.
@@ 4.2 Modeling framework
$\phi^{\mathrm{a}}_{B^{+}}(\mathbf{b}_{2})=\int_{\Gamma^{\mathrm{a}}_{\mathbf{b}_{2}}}d\varphi_{B^{+}}\,$, (18)
where the differential one-form $d\varphi_{B^{+}}$ is analogous to $d\varphi_{W^{+}}$. This simplified framework is mathematically convenient because $B$ now has the same role as $W$. In particular, we call $\mathbf{b}_{2}$ an absolute phase singularity if the phase is undefined somewhere along $\Gamma^{\mathrm{a}}_{\mathbf{b}_{2}}$, at an intermediate time. The domain of $\phi^{\mathrm{a}}_{B^{+}}$ is denoted $B^{\mathrm{a}}$. As for $\phi^{\mathrm{a}}$, we will adopt the convex combination absolute path (15). For instance, if $\mathbf{b}$ encodes scatterer positions, the scatterers will move at a uniform rate from $\mathbf{b}_{1}$ to $\mathbf{b}_{2}$.
@@ 4.3.1 Point scatterer
For a point scatterer, the absolute phase defined through (17) is proportional to the difference in optical path length, as expected [bamler98]. Assuming a monostatic configuration, zero spatial baseline, the plane wave approximation and no atmospheric or other phase screens, the single-look coherence
$\gamma_{\mathrm{point}}(t)=\exp\left(\mathsf{j}2k\left[R(t)-R(t_{1})\right]\right)\,$, (19)
where $k$ is the scalar wavenumber, $R$ the range (line-of-sight distance to the scatterer), and the primary acquisition at $t_{1}$ is fixed. Multilook observations of a non-decorrelating, uniformly moving surface are also described by (19). Previously [bamler98, rosen00], the absolute phase of such scatterers was taken as
$\phi^{\mathrm{a}}_{\mathrm{point}}(t_{2};t_{1})=2k\left[R(t_{2})-R(t_{1})\right]\,.$ (20)
@@
$d\varphi_{t}(\mathrm{d}t)=2k\frac{\mathrm{d}R}{\mathrm{d}t}\,\mathrm{d}t=2k\,\mathrm{d}R\,$ (21)
is an exact form, so the integral only depends on the range difference $R(t_{2})-R(t_{1})$ and not on the path $R(t)$.
@@ 4.3.2 Three surface segments
The three surface segments from the introduction differ from a single surface or point scatterer in that a cyclic change can induce a non-zero absolute phase. To model the system, let the state $\mathbf{b}=[b_{1},b_{2},b_{3}]$ encode the position of the three segments, which are not resolved in the observations. Assuming the returns from different segments are uncorrelated and of equal strength [zwieback24], let $b_{i}(s)$ be the position of surface $i$ at dimensionless time $s$, nondimensionalized such that the modeled population coherence is
$\gamma(s)=\frac{1}{3}\sum_{i=1}^{3}\exp^{-\mathsf{j}\left(b_{i}(s)-b_{i}(0)\right)}\,.$ (22)
@@
Each segment starts out at $b_{i}(0)=0$ for the fixed primary acquisition at $s=0$ and acts like a non-decorrelating point-like scatterer from (19). Fig. 3a shows a loop that corresponds to the cyclic change in surface positions from Fig. 1. As all segments return to their initial position at $s=1$, a nonzero value of the absolute phase $\phi^{\mathrm{a}}(1)=-2\pi$ is obtained.
@@
Figure 3: Differential InSAR model of three surface segments with positions $\mathbf{b}=[b_{1},b_{2},b_{3}]$. a) The purple loop with basepoint $\mathbf{b}_{1}=[0,0,0]$ (white cross) in parameter space describes the temporal evolution from Fig. 1, it encircles a phase singularity (black line) in the plane $b_{1}=0$. b–d) Wrapped phase with white contours indicating vanishing real and imaginary part (b), coherence (c) and absolute phase in the modeling framework of Sec. 4.2 (d) shown in the plane $b_{1}=0$ for fixed primary $\mathbf{b}_{1}$, along with phase singularities (black dots) and absolute phase singularities (gray lines).
The nonzero absolute phase for the cyclic change demonstrates that the absolute phase is not proportional to a range difference, which may seem counterintuitive. In the case of the three surface segments, an obvious way to measure the range difference with respect to $\mathbf{b}_{1}=[0,0,0]$ is the average of $\mathbf{b}=[b_{1},b_{2},b_{3}]$. A weighted average would be another option. In general, let us denote by $\Delta r(\mathbf{b})$ any measure of the range difference that vanishes whenever $\mathbf{b}=\mathbf{b}_{1}\equiv[0,0,0]$ and that is continuous in $\mathbf{b}$. To combat the counter-intuitive nature of the absolute phase defined through (17), couldn’t we define an alternative absolute phase that is proportional to $\Delta r$? The answer is no, as long we require that absolute phase to be congruent with the wrapped phase (making it a phase-like quantity), to depend continuously on $\mathbf{b}(s)$ (as $\Delta r$ does), and to equal zero when the surface positions do not change (as $\Delta r$ does). For any continuous temporal trajectory $\mathbf{b}(s)$, these three requirements entail (Sec. 4.1) the observational definition of the absolute phase in (17). The non-zero value $\phi^{\mathrm{a}}(1)=-2\pi$ for the cyclic change implies that the absolute phase cannot be proportional to any $\Delta r$. To develop insight into the origin of the loss of proportionality to a range difference, we next consider the phase in parameter space.
@@
The nonzero absolute phase around the loop relates to the loop enclosing a phase singularity in parameter space, one of the two dark lines in Fig. 3a). In the plane $b_{1}=0$, the wrapped phase $\phi(\mathbf{b})$ for fixed primary $\mathbf{b}_{1}=0$ in Fig. 3b spirals around the phase singularity, such that the unwrapped phase around the loop gives the absolute phase $\phi^{\mathrm{a}}(1)=-2\pi$. As $\rho>0$ along the loop (Fig. 3c), this absolute phase is well defined.
@@
The phase singularities induce absolute phase singularities in the simplified modeling framework of Sec. 4.2. While the loop does not fit into this framework, let us impose the framework by restricting trajectories to the line segment from $\mathbf{b}_{1}=0$ to secondary $\mathbf{b}$. The absolute phase in parameter space, $\phi^{\mathrm{a}}_{B}$ is shown in Fig. 3d). The absolute phase singularities $A\in B$ in the plane $b_{1}=0$ are rays extending from each phase singularity, across which the absolute phase jumps by $2\pi$. They leave no trace in the wrapped phase. In Sec. 6.1, it will be shown that the discontinuous jumps arise from the path-dependence of the phase one-form illustrated by the loop encircling the phase singularity.